# Tracezip: Efficient Distributed Tracing via Trace Compression

### v1

@@ Abstract.
Distributed tracing serves as a fundamental building block in the monitoring and testing of cloud service systems. To reduce computational and storage overheads, the de facto practice is to capture fewer traces via sampling. However, existing work faces a trade-off between the completeness of tracing and system overhead. On one hand, head-based sampling indiscriminately selects requests to trace when they enter the system, which may miss critical events. On the other hand, tail-based sampling traces all requests and selectively persist the edge-case traces, which entails the overheads related to trace collection and ingestion. Taking a different path, in this paper we propose Tracezip to enhance the efficiency of distributed tracing via trace compression. Our key insight is that there exists significant redundancy among traces, which results in repetitive transmission of identical data between the services and backend. We design a new data structure named Span Retrieval Tree (SRT) that continuously encapsulates such redundancy at the service side and transforms trace spans into a lightweight form. At the backend, the full traces can be seamlessly reconstructed by retrieving the common data already delivered by previous spans. Tracezip includes a series of strategies to optimize the structure of SRT and a differential update mechanism to efficiently synchronize SRT between services and backend. Our evaluation on microservices benchmarks, popular cloud service systems, and production trace data demonstrate that Tracezip can achieve substantial performance gains in trace collection, with negligible overhead. We have implemented Tracezip inside OpenTelemetry Collector, making it compatible with existing tracing APIs.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The ACM SIGSOFT International Symposium on Software Testing and Analysis; June 25–28, 2025; Trondheim, Norwayisbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organization Cloud computingccs: Computer systems organization Reliabilityccs: Computer systems organization Maintainability and maintenanceccs: Computer systems organization Redundancyccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
In modern cloud systems, the adoption of loosely coupled designs for applications and services has marked a significant paradigm shift in software architecture. While such modular design brings the benefits of scalability and operational flexibility, it also complicates different aspects of software development and maintenance, including testing, debugging and diagnosing. This can be largely attributed to the cascading effect of failures (Dang et al., 2019; Chen et al., 2020), i.e., a single failure in one service can quickly propagate to other interconnected services. As such, distributed tracing has emerged as an essential reliability management solution in cloud systems. This is primarily due to its ability to provide a comprehensive and granular view of the interactions between services, allowing for the precise identification of where and how failures occur and spread.
In cloud service systems, unusual and edge-case system behaviors, such as tail latency, are rare by definition. To achieve high coverage of outlier system events, it is necessary to trace all requests. In production environments, this may result in substantial trace data, incurring significant overhead and costs related to trace generation, collection, and ingestion. To mitigate this problem, various trace sampling techniques have been proposed, which can be categorized into two main types, i.e., head-based sampling (Zhang et al., 2023) and tail-based sampling (Chen et al., 2024; Huang et al., 2021; He et al., 2023; Zhang et al., 2023). Head-based sampling uniformly collects traces at random based on a small sampling rate (e.g., 1% (Zhang et al., 2023; Las-Casas et al., 2019; Kaldor et al., 2017)). The sampling decision is made before request execution, and only the sampled requests will be traced. On the other hand, tail-based sampling captures traces for all requests, and decides whether to retain a trace based on the execution details such as latency and HTTP status code. In this regard, many machine learning techniques have been applied to automatically select informative and uncommon traces. However, given the inherent unpredictability of the true value of traces, trace sampling struggles to capture the full spectrum of critical information necessary for effective software testing and failure diagnosis.
@@
To pursue more effective compression, Tracezip harnesses the global redundancy inherent in the structures of trace spans. Specifically, we design a new data structure, Span Retrieval Tree (SRT), based on the principles of prefix trees, which is able to continuously extract the set of key-value pairs commonly shared across spans. The SRT, when synchronized between services and backend, serves as a reference mechanism to retrieve the identical data that have been previously transmitted by other spans. To manage computational and space complexity, Tracezip constantly restructures SRT into its most compact form and employs mapping techniques to further reduce its size. We also propose a differential update mechanism to effectively synchronize SRT between services and backend. Tracezip is orthogonal to trace sampling methods and can work with log compression techniques once trace data have been efficiently transmitted to the backend.
We have implemented Tracezip inside OpenTelemetry Collector (OpenTelemetry, 2024e), one of the most popular tracing frameworks, making it compatible with existing tracing APIs. We deploy Tracezip to collect traces for an open-source microservices benchmark and six application backend components in cloud environments, i.e., gRPC, Apache Kafka, Servlet, MySQL, Redis, MongoDB. We also evaluate the compression performance of Tracezip on Alibaba production traces. The experimental results show that Tracezip offers around 10$\sim$45performance gain when working in conjunction with traditional compression schemes, i.e., gzip, bzip2, lzma. Moreover, Tracezip demonstrates negligible space overhead (i.e., several megabytes) and high efficiency.
@@
  • We propose Tracezip, the first online trace compression system by leveraging the inherent redundancy in trace span data. This information is captured by a new data structure Span Retrieval Tree (SRT). By sharing SRT between services and tracing backend, we can eliminate the redundancy associated with the repetitive transmission of identical data across multiple spans.
@@ 2.1. Distributed Tracing
Figure 1 illustrates a typical procedure (Zhang et al., 2023; OpenTelemetry, 2024a) of tracking requests in modern distributed tracing frameworks, such as Jaeger (Jaeger, 2024) and Zipkin (Zipkin, 2024). Upon the arrival of a new request to the application, the tracing framework assigns it a unique trace_id (➀). However, not all requests will actually be traced, which is indicated by a flag, sampled. Both trace_id and sampled will then be propagated along the request at the application level, which is important for the completeness and coherence of a trace. If sampled is set, each component (e.g., a microservice instance) that handles the request will generate trace data (➁), i.e., a span, using the tracing framework’s client library (e.g., OpenTelemetry). The places that emit spans are called a tracepoint, and there could be multiple tracepoints serving different requests or different operations within a single request. The framework’s client library then enqueues, serializes, and transmits trace spans (➂) to its centralized backend collection infrastructure, or simply backend. The backend is responsible for continuously receiving (➃), processing (➄), and storing (➅) trace data. Based on the trace_id, parent_span_id, and span_id, the backend can assemble spans that were dispersed across different components into a single coherent trace.
Given the details provided by these spans, operators can closely monitor and understand the impact that one component may have on the others. This makes trace data particularly useful for troubleshooting cross-component problems in large distributed systems. However, traces can be produced at high volume, incurring significant network, compute, and storage costs. In production scenarios, Google is estimated to generate approximately 1,000 TB of raw traces on a daily basis. Netflix needs to manage more than 2 billions of daily requests. To mitigate overheads, existing tracing frameworks often apply head-based sampling to trace only a small fraction of requests by setting the sampled flag (➀). This will inevitably increases the risk of overlooking system edge-case behaviors. On the other hand, tail-based sampling utilizes a filtering strategy, i.e., only persisting traces that exhibit outliers symptoms (➅), e.g., high tail latency, error codes. However, tail-based sampling entails enormous costs, as it must trace all requests and ingest the trace data to make sampling decisions.
@@ 2.2. A Study of the Redundancy in Trace Data
The studied systems include an open-source microservices benchmark, i.e., Train Ticket, which is popular in cloud-related research fields, and six application backend components that are selected for their widespread use in modern cloud systems, i.e., gRPC (Client and Server), Apache Kafka (Producer and Consumer), Servlet, MySQL, Redis, and MongoDB. They cover a diverse functionalities including message queuing, HTTP communication, remote procedure calls, database management, etc. To collect their traces, we leverage the zero-code instrumentation capabiltiy of OpenTelemetry (OpenTelemetry, 2024c). It allows the collection of observability data, i.e., logs, metrics, and traces, for applications without the need to modify the source code. This is achieved by using libraries, plugins, or agent to instrument the libraries used by applications. It supports many programming languages (e.g., Java, Go, Python) and a wide range of popular libraries and frameworks, including requests and responses, database calls, message queue calls. By injecting typical workloads to the studied systems (Section 5.1.1), we collect more than 40GB of trace data in total.
We quantify trace redundancy by measuring the proportion of duplicate key-value (KV) pairs generated by each service. As a KV pair represents the fundamental unit of information within a trace, assessing its repetition enables us to gauge the degree of information overlap. Specifically, for each KV pair, we first count the number of its occurrence and then calculate its ratio over the total number of KV pairs in the dataset. For example, in a trace dataset containing 100 KV pairs, if two specific pairs appear once and ten times respectively, their redundancy ratios would be 1% and 10%, respectively. We calculate the fraction of KV pairs within each service which occur less than 1,000 times, as well as those exceeding this threshold, as shown in Figure 2. We made two observations:
@@
Traces are highly redundant. Based on the figures, we can see that for different microservices in Train Ticket, around 70% of the total KV pairs are highly repetitive. Similar situations can be found in the application components. The only exception is Apache Kafka, whose KV pairs tend to be unique. The reason behind is that Kafka’s traces include the data from its message queues, increasing the randomness of its KV pairs. We will discuss the impact during system evaluation (Section 5). The results indicate that a significant portion of the trace data is characterized by redundant information. An important reason is that services often engage in standard interactions and perform routine operations that generate trace data with similar patterns. By capturing the redundancy upon the generation of spans at the service side (➀), we can preemptively eliminate the transmission of repetitive data that already exist in the backend. The backend can easily reconstruct the complete spans based on their redundancy patterns (➅).
There exists structural redundancy among attributes. We also observe that there exists certain redundancy at span level. OpenTelemetry’s semantic conventions (OpenTelemetry, 2024d) offer standardized guidelines for naming common attributes across different kinds of operations and data. This is essential for maintaining the uniformity and compatibility of naming scheme across languages, libraries, and platforms. In the naming scheme, attribute names are organized into hierarchical namespaces to indicate their context or category, e.g., network.local.address, network.local.port, network.peer.address. We can see that they share some common words. The attribute values also exhibit similar commonality, e.g., Java Exception has java.io.IIOException, java.io.EOFException. By removing such fine-grained redundancy, we can further reduce the overheads of span transmission.
@@ 3.1. Overview
In this section, we present the design of Tracezip. The system architecture is illustrated in Figure 3, in which we add a compression module (➆) to the service side and a corresponding decompression module (➇) to the entry of the backend trace collectors. In the compression module, we maintain two data structures, namely, a Span Retrieval Tree (SRT) and a dictionary, to constantly capture the redundancy across the spans. Upon the generation of a span at the tracepoint, it undergoes compression utilizing the above structures. If the span carries a new redundancy pattern, it will be seamlessly integrated into the SRT and dictionary. This integration is crucial as it enriches the structures, thereby facilitating the compression for subsequent spans. We accelerate the above process by employing a combination of mapping and hashing techniques. At the decompression module, the spans are restored to their original form by referring to the SRT and the dictionary. To ensure a consistent and reliable data transmission, it is imperative that these data structures are accurately synchronized between the service and backend sides. To achieve this, we develop a differential update mechanism (➈). This mechanism is designed to precisely pinpoint and propagate only the incremental changes in the data structures, ensuring an efficient synchronization process that minimizes overhead while maximizing data consistency.
@@ 3.2. Span Format Conventions
To compress spans by leveraging their recurring patterns, we first stipulate the format of a span. For simplicity and readibility, we assume that a span adheres to the standard JSON data format that defines it as a structured set of key-value pairs. The key is a string, while the value can be primitive types (strings, numbers, booleans, and null) or two structured types (nested key-value pairs and arrays). This aligns with the format specifications used in many tracing frameworks and tools, e.g., OpenTelemetry (OpenTelemetry, 2024b), Jaeger (Jaeger, 2024), Zipkin (Zipkin, 2024). Typical fields (keys) of a span include: Name (a human-readable string representing the operation done), Parent Span ID (the span that caused the creation of this span, empty for root spans), Start and End Timestamps (the start and end time of the span), Span Context (the context of the span including the trace ID, the span ID, etc.), Attributes (key-value pairs representing additional information about the span), Span Events (structured log messages/annotations on a span), etc. It is important to note that our proposed algorithm is not restricted to JSON or any particular serialization format. For example, Tracezip can work effectively when Protobuf (Protocol Buffers) (Buffers, 2024) is used for trace data serialization. With Protobuf’s powerful deserialization capabilities, we can leverage its reflection-like APIs or direct-access methods to dynamically access the fields and values of spans. Additionally, Protobuf is designed to be backward and forward compatible, allowing us to modify the message definition by adding or removing fields while maintaining compatibility with older data. Such operations are essential for reducing trace redundancy, e.g., removing span elements that are deemed repetitive.

### v2

@@ Abstract.
Distributed tracing serves as a fundamental building block in the monitoring and testing of cloud service systems. To reduce computational and storage overheads, the de facto practice is to capture fewer traces via sampling. However, existing work faces a trade-off between the completeness of tracing and system overhead. On one hand, head-based sampling indiscriminately selects requests to trace when they enter the system, which may miss critical events. On the other hand, tail-based sampling first captures all requests and then selectively persists the edge-case traces, which entails the overheads related to trace collection and ingestion. Taking a different path, we propose Tracezip in this paper to enhance the efficiency of distributed tracing via trace compression. Our key insight is that there exists significant redundancy among traces, which results in repetitive transmission of identical data between services and the backend. We design a new data structure named Span Retrieval Tree (SRT) that continuously encapsulates such redundancy at the service side and transforms trace spans into a lightweight form. At the backend, the complete traces can be seamlessly reconstructed by retrieving the common data that are already delivered by previous spans. Tracezip includes a series of strategies to optimize the structure of SRT and a differential update mechanism to efficiently synchronize SRT between services and the backend. Our evaluation on microservices benchmarks, popular cloud service systems, and production trace data demonstrates that Tracezip can achieve substantial performance gains in trace collection with negligible overhead. We have implemented Tracezip inside the OpenTelemetry Collector, making it compatible with existing tracing APIs.
@@
copyright: acmlicensedjournal: PACMSEjournalyear: 2025journalvolume: 2journalnumber: ISSTAarticle: ISSTA019publicationmonth: 7doi: 10.1145/3728888ccs: Computer systems organization Cloud computingccs: Computer systems organization Reliabilityccs: Computer systems organization Maintainability and maintenanceccs: Computer systems organization Redundancyccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
In modern cloud systems, the adoption of loosely coupled designs for applications and services has marked a significant paradigm shift in software architecture. While such modular design brings the benefits of scalability and operational flexibility, it also complicates different aspects of software development and maintenance, including testing, debugging and diagnosing. This can be largely attributed to the cascading effect of failures (Dang et al., 2019; Chen et al., 2020), i.e., a single failure in one service can quickly propagate to other interconnected services and components. As such, distributed tracing has emerged as an essential reliability management solution in cloud systems. This is primarily due to its ability to provide a comprehensive and granular view of the interactions between services, allowing for the precise identification of where and how failures occur and spread.
In cloud service systems, unusual and edge-case system behaviors, such as tail latency and resource contention, are rare by definition. To achieve high coverage of outlier system events, it is necessary to trace all requests. In production environments, this may result in substantial trace data, incurring significant overhead and costs related to trace generation, collection, and ingestion. To mitigate this problem, various trace sampling techniques have been proposed, which can be categorized into two main types, i.e., head-based sampling (Zhang et al., 2023) and tail-based sampling (Chen et al., 2024; Huang et al., 2021; He et al., 2023; Zhang et al., 2023). Head-based sampling uniformly collects traces at random based on a small sampling rate (e.g., 1% (Zhang et al., 2023; Las-Casas et al., 2019; Kaldor et al., 2017)). The sampling decision is made before request execution, and only the sampled requests will be traced. On the other hand, tail-based sampling captures traces for all requests, and decides whether to retain a trace based on the execution details such as latency and HTTP status code. In this regard, many machine learning techniques have been applied to automatically select informative and uncommon traces. However, given the inherent unpredictability of the true value of traces, trace sampling struggles to capture the full spectrum of critical information necessary for effective software testing and failure diagnosis.
@@
To pursue more effective compression, Tracezip harnesses the global redundancy inherent in the structures of trace spans. Specifically, we design a new data structure, Span Retrieval Tree (SRT), based on the principles of prefix trees, which is able to continuously extract the set of key-value pairs commonly shared across spans. The SRT, when synchronized between services and the backend, serves as a reference mechanism to retrieve the identical data that have been previously transmitted by other spans. To manage computational and space complexity, Tracezip constantly restructures SRT into its most compact form and employs mapping techniques to further reduce its size. We also propose a differential update mechanism to effectively synchronize SRT between services and the backend. Tracezip is orthogonal to trace sampling methods and can work with log compression techniques once trace data have been efficiently transmitted to the backend.
We have implemented Tracezip inside OpenTelemetry Collector (OpenTelemetry, 2024e), one of the most popular tracing frameworks, making it compatible with existing tracing APIs. We deploy Tracezip to collect traces for an open-source microservices benchmark (i.e., Train Ticket (Zhou et al., 2021)) and six application backend components in cloud environments (i.e., gRPC, Apache Kafka, Servlet, MySQL, Redis, MongoDB). We also evaluate the compression performance of Tracezip on Alibaba production traces (Luo et al., 2021; Wang et al., 2022). The experimental results show that Tracezip offers around 10$\sim$45performance gain when working in conjunction with traditional compression schemes, i.e., gzip, bzip2, lzma. Moreover, Tracezip demonstrates negligible space overhead (i.e., several megabytes) and high efficiency.
@@
  • We propose Tracezip, the first online trace compression system by leveraging the inherent redundancy in trace span data. This information is captured by a new data structure we developed, the Span Retrieval Tree (SRT). By sharing SRT between services and the tracing backend, we can eliminate the redundancy associated with the repetitive transmission of identical data across multiple spans.
@@ 2.1. Distributed Tracing
Figure 1 illustrates a typical procedure of tracking requests (Zhang et al., 2023; OpenTelemetry, 2024a) in modern distributed tracing frameworks, such as Jaeger (Jaeger, 2024) and Zipkin (Zipkin, 2024). Upon the arrival of a new request to the application, the tracing framework assigns it a unique trace_id (➀). However, not all requests will actually be traced, which is indicated by a flag, sampled. Both trace_id and sampled will then be propagated along the request at the application level, which is important for the completeness and coherence of a trace. If sampled is set, each component (e.g., a microservice instance) that handles the request will generate trace data (➁), i.e., a span, using the tracing framework’s client library (e.g., OpenTelemetry). The places that emit spans are called a tracepoint, and there could be multiple tracepoints serving different requests or different operations within a single request. The framework’s client library then enqueues, serializes, and transmits trace spans (➂) to its centralized backend collection infrastructure, or simply backend. The backend is responsible for continuously receiving (➃), processing (➄), and storing (➅) trace data. Based on the trace_id, parent_span_id, and span_id, the backend can assemble spans that were dispersed across different components into a single coherent trace.
Given the details provided by these spans, operators can closely monitor and understand the impact that one component may have on the others. This makes trace data particularly useful for troubleshooting cross-component problems in large distributed systems. However, traces can be produced at high volume, incurring significant network, compute, and storage costs. In production scenarios, Google is estimated to generate approximately 1,000 TB of raw traces on a daily basis. Netflix needs to manage more than 2 billions of daily requests. To mitigate overheads, existing tracing frameworks often apply head-based sampling to trace only a small fraction of requests by randomly setting the sampled flag (➀). This will inevitably increases the risk of overlooking system edge-case behaviors. On the other hand, tail-based sampling utilizes a filtering strategy, i.e., only persisting traces that exhibit outliers symptoms (➅), e.g., high tail latency, error codes. However, tail-based sampling entails enormous costs, as it must trace all requests and ingest the trace data to make sampling decisions.
@@ 2.2. A Study of the Redundancy in Trace Data
The studied systems include an open-source microservices benchmark, i.e., Train Ticket (Zhou et al., 2021), which is popular in cloud-related research fields, and six application backend components that are selected for their widespread use in modern cloud systems, i.e., gRPC (Client and Server), Apache Kafka (Producer and Consumer), Servlet, MySQL, Redis, and MongoDB. They cover a diverse functionalities including message queuing, HTTP communication, remote procedure calls, database management, etc. To collect their traces, we leverage the zero-code instrumentation capabiltiy of OpenTelemetry (OpenTelemetry, 2024c). It allows the collection of observability data, i.e., logs, metrics, and traces, for applications without the need to modify the source code. This is achieved by using libraries, plugins, or agent to instrument the libraries used by applications. It supports many programming languages (e.g., Java, Go, Python) and a wide range of popular libraries and frameworks, including requests and responses, database calls, message queue calls. By injecting typical workloads to the studied systems (Section 5.1.1), we collect more than 40GB of trace data in total.
We quantify trace redundancy by measuring the proportion of duplicate key-value (KV) pairs generated by each service. As a KV pair represents the fundamental unit of information within a trace, assessing its repetition enables us to gauge the degree of information overlap. Specifically, for each KV pair, we first count the number of its occurrence and then calculate its ratio over the total number of KV pairs in the dataset. For example, in a trace dataset containing 100 KV pairs, if two specific pairs appear once and ten times respectively, their redundancy ratios would be 1% and 10%, respectively. We calculate the fraction of KV pairs within each service which occur less than 1,000 times, as well as those exceeding this threshold, as shown in Figure 2. We make the following two important observations.
@@
Traces are highly redundant. Based on the figures, we can see that for different microservices in Train Ticket, around 70% of the total KV pairs are highly repetitive. Similar situations can be found in the application components. The only exception is Apache Kafka, whose KV pairs tend to be unique. The reason behind is that Kafka’s traces include the data from its message queues, increasing the randomness of its KV pairs. We will discuss its impact on compression during system evaluation (Section 5). The results indicate that a significant portion of the trace data is characterized by redundant information. An important reason is that services often engage in standard interactions and perform routine operations that generate trace data with similar patterns. By capturing the redundancy upon the generation of spans at the service side (➀), we can preemptively eliminate the transmission of repetitive data that already exist in the backend. The backend can easily reconstruct the complete spans based on their redundancy patterns (➅).
There exists structural redundancy among attributes. We also observe that there exists certain redundancy at span level. OpenTelemetry’s semantic conventions (OpenTelemetry, 2024d) offer standardized guidelines for naming common attributes across different kinds of operations and data. This is essential for maintaining the uniformity and compatibility of naming scheme across languages, libraries, and platforms. In the naming scheme, attribute keys are organized into hierarchical namespaces to indicate their context or category, e.g., network.local.address, network.local.port, network.peer.address. We can see that they share some common words. The attribute values also exhibit similar commonality, e.g., Java Exception has java.io.IIOException, java.io.EOFException. By eliminating such fine-grained redundancies, we can further reduce the overhead associated with span transmission.
@@ 3.1. Overview
In this section, we present the design of Tracezip. The system architecture is illustrated in Figure 3, in which we add a compression module (➆) to the service side and a corresponding decompression module (➇) to the entry of the backend trace collectors. In the compression module, we maintain two data structures, namely, a Span Retrieval Tree (SRT) and a dictionary, to constantly capture the redundancy across the spans. Upon the generation of a span at the tracepoint, it undergoes compression utilizing the above data structures. If the span carries a new redundancy pattern, it will be seamlessly integrated into the SRT and dictionary. This integration is crucial as it enriches the structures, thereby facilitating the compression for subsequent spans. We accelerate the above process by employing a combination of mapping and hashing techniques. At the decompression module, the spans are restored to their original form by referring to the SRT and the dictionary. To ensure a consistent and reliable data transmission, it is imperative that these data structures are accurately synchronized between the service and backend sides. To achieve this, we develop a differential update mechanism (➈). This mechanism is designed to precisely pinpoint and propagate only the incremental changes in the data structures, ensuring an efficient synchronization process that minimizes overhead while maximizing data consistency.
@@ 3.2. Span Format Conventions
To compress spans by leveraging their recurring patterns, we first stipulate the format of a span. For simplicity and readibility, we assume that a span adheres to the standard JSON data format that defines it as a structured set of key-value pairs. The key is a string, while the value can be either primitive types (strings, numbers, booleans, and null) or two structured types (nested key-value pairs and arrays). This aligns with the format specifications used in many tracing frameworks and tools, e.g., OpenTelemetry (OpenTelemetry, 2024b), Jaeger (Jaeger, 2024), Zipkin (Zipkin, 2024). Typical fields (keys) of a span include: Name (a human-readable string representing the operation done), Parent Span ID (the span that caused the creation of this span, empty for root spans), Start and End Timestamps (the start and end time of the span), Span Context (the context of the span including the trace ID, the span ID, etc.), Attributes (key-value pairs representing additional information about the span), Span Events (structured log messages/annotations on a span), etc. It is important to note that our proposed algorithm is not restricted to JSON or any particular serialization format. For example, Tracezip can work effectively when Protobuf (Protocol Buffers) (Buffers, 2024) is used for trace data serialization. With Protobuf’s powerful deserialization capabilities, we can leverage its reflection-like APIs or direct-access methods to dynamically access the fields and values of spans. Additionally, Protobuf is designed to be backward and forward compatible, allowing us to modify the message definition by adding or removing fields while maintaining compatibility with older data. Such operations are essential for reducing trace redundancy, e.g., removing span elements that are deemed repetitive.